# Full-Duplex-Bench: A Benchmark to Evaluate Full-Duplex Spoken Dialogue Models on Turn-taking Capabilities

### v2

@@ Abstract
Spoken dialogue modeling poses challenges beyond text-based language modeling, requiring real-time interaction, turn-taking, and backchanneling. While most Spoken Dialogue Models (SDMs) operate in half-duplex mode—processing one turn at a time—emerging full-duplex SDMs can listen and speak simultaneously, enabling more natural conversations. However, current evaluations remain limited, focusing mainly on turn-based metrics or coarse corpus-level analyses. To address this, we introduce Full-Duplex-Bench [^1] Data, codebase, and demo samples will be released here https://github.com/DanielLin94144/Full-Duplex-Bench., a benchmark that systematically evaluates key interactive behaviors: pause handling, backchanneling, turn-taking, and interruption management. Our framework uses automatic metrics for consistent, reproducible assessment and provides a fair, fast evaluation setup. By releasing our benchmark and code, we aim to advance spoken dialogue modeling and foster the development of more natural and engaging SDMs.
@@ I Introduction
TABLE I: Overview of full-duplex speech language models. “—” indicates properties that are not publicly available or unspecified. “E2E” denotes end-to-end speech modeling (without relying on text), “#ch” indicates the number of speech input channels, “Interrupt” refers to interruption handling, “BC” represents backchanneling capability, and “S2S Release” shows whether the complete speech-to-speech pipeline is publicly released (some models only release partial checkpoints without this functionality).
@@
In this paper, we address the need by introducing Full-Duplex-Bench, the first benchmark designed to systematically evaluate key turn-taking behaviors in full-duplex spoken dialogue models (SDMs), as illustrated in Figure 1. Our framework targets four critical aspects of real-time interaction: pause handling, backchanneling, turn-taking, and user interruption management, using automatic metrics to enable consistent and reliable comparisons. This benchmark represents an important initial step toward evaluating the full-duplex turn-taking capabilities of SDMs. Moreover, the benchmark can be executed efficiently and automatically, enabling rapid evaluation across models. By applying our benchmark to existing systems, we gain deeper insights into how they manage interactive dynamics in real time. We will publicly release the benchmark and associated codebase to encourage further research and the development of more engaging and conversational SDMs.
@@ III-B Evaluation Dimensions
TABLE II: The number of samples for each dataset and task in Full-Duplex-Bench.
@@ III-C Data Curation
For the pause handling task, we first apply voice activity detection to identify segments with no overlapping speech, ensuring that only one speaker is active and dominant. We then select turns that contain an internal pause lasting between 0.4 and 1 second. Additionally, we filter out turns that are too short (under 5 seconds) and exclude cases where the speech before or after the pause consists of backchannel responses.
For the smooth turn-taking task, we compute the gap between each pair of consecutive turns and retain only those where all gaps are less than 0.4 seconds. We ensure that there is no overlapping speech, each turn is longer than 4 seconds, and that the speech is not backchanneling. To allow the evaluated model sufficient time to respond, we append 5 seconds of silence at the end of the input stream. After automatic filtering, we manually review the audio samples to ensure quality and consistency.
@@
Synthetic Data (User Interruption, Pause Handling): To address the scarcity of user interruptions in existing public datasets, we generated synthetic dialogues using GPT-4o [45], which include both contextual turns and interruption turns. Text-to-Speech synthesis was performed using ChatTTS [^6] https://github.com/2noise/ChatTTS, with 10 different speaker voices randomly assigned across samples to enhance diversity. For each input stream fed to the evaluated model, the interrupting speech is played 7 seconds after the preceding utterance. Additionally, we append 15 seconds of silence after the interruption to allow the model time to respond.
@@ IV Models Under Evaluation
Since many models do not release complete speech-to-speech checkpoints (see Table I, column “S2S Release”) [^7] By open-sourcing our data and metrics, we encourage the broader community, including those with closed-source models, to evaluate their systems using our benchmark, currently we evaluate models with a publicly available speech-to-speech model and inference pipeline. We also include the recently released Gemini Live API as a representative commercial model due to its user-friendly interface.
  • dGSLM [26]: A textless speech-to-speech model that generates natural conversations directly from audio. It uses (1) a HuBERT+k-means encoder [46], (2) a dual-tower Transformer for two-channel dialogue modeling, and (3) a HiFi-GAN decoder. Trained on 2,000 hours of phone calls [47], it captures both linguistic and paralinguistic cues. Originally non-interactive, we adapt it for live interaction using the official implementation [^8] https://github.com/facebookresearch/fairseq/tree/main/examples/textless_nlp/dgslm with our modifications in the Appendix.
  • Moshi [7]: A real-time speech-to-speech system combining a 7B LLM (Helium) and neural codec (Mimi) via residual vector quantization. It features an “Inner Monologue” step to improve fluency and supports overlapping speech and interruptions through a multi-stream architecture. We use the official implementation [^9] https://github.com/kyutai-labs/moshi; setup details are in the Appendix.
@@ VI Conclusion
In this paper, we introduced Full-Duplex-Bench, a benchmark designed to evaluate critical aspects of full-duplex spoken dialogue models. Our framework targets key interaction dimensions—pause handling, backchanneling, smooth turn-taking, and user interruption management—addressing limitations of existing benchmarks that primarily focus on half-duplex settings or coarse corpus-level metrics. To ensure systematic and reproducible evaluation, we propose automatic metrics tailored to real-time interaction. Experiments on open-source full-duplex models reveal distinct model features and highlight areas for improvement. We hope Full-Duplex-Bench provides a practical foundation for evaluating full-duplex spoken dialogue systems. By releasing our data and behavior-specific metrics, we aim to enable fair comparison and advance research toward more natural and responsive human-computer interaction.

### v3

@@ Abstract
Spoken dialogue modeling poses challenges beyond text-based language modeling, requiring real-time interaction, turn-taking, and backchanneling. While most Spoken Dialogue Models (SDMs) operate in half-duplex mode—processing one turn at a time—emerging full-duplex SDMs can listen and speak simultaneously, enabling more natural conversations. However, current evaluations remain limited, focusing mainly on turn-based metrics or coarse corpus-level analyses. To address this, we introduce Full-Duplex-Bench, a benchmark that systematically evaluates key interactive behaviors: pause handling, backchanneling, turn-taking, and interruption management. Our framework uses automatic metrics for consistent, reproducible assessment and provides a fair, fast evaluation setup. By releasing our benchmark and code, we aim to advance spoken dialogue modeling and foster the development of more natural and engaging SDMs.
@@ I Introduction
TABLE I: Overview of full-duplex speech language models. “—” indicates properties that are not publicly available or unspecified. “E2E” denotes end-to-end speech modeling (without relying on text), “#ch” indicates the number of speech input channels, “Interrupt” refers to interruption handling, “BC” represents backchanneling capability, and “S2S Release” shows whether the complete speech-to-speech pipeline is publicly released.
@@
In this paper, we present Full-Duplex-Bench, the first scenario-driven benchmark for systematically evaluating key turn-taking behaviors in human-to-machine full-duplex spoken dialogue models (SDMs), as illustrated in Figure 1. The benchmark evaluates models under realistic conversational events such as backchannels and user interruptions, focusing on four critical aspects of real-time interaction: pause handling, backchanneling, turn-taking, and user interruption management. Our unified framework offers objective behavior detection, rapid diagnostic metrics, and an open, reproducible toolkit, enabling consistent and reliable cross-model comparisons. The metrics are intentionally descriptive rather than prescriptive, allowing developers to prioritize behaviors according to specific application requirements. Designed for efficiency and full automation, the benchmark supports rapid large-scale evaluation and provides deeper insights into how existing systems manage interactive dynamics in real time. We publicly release the benchmark and codebase to facilitate further research and the development of more engaging and conversational SDMs [^1] Data and code are released in https://github.com/DanielLin94144/Full-Duplex-Bench..
@@ III-B Evaluation Dimensions
TABLE II: The number of samples for each dataset and task.
@@ III-C Data Curation
For the pause handling task, we first apply voice activity detection to identify segments with no overlapping speech, ensuring that only one speaker is active. We then select turns containing an internal pause between 0.4 and 1.0 seconds. This range is informed by prior research showing that pauses of around one second are perceived as a natural upper bound in English conversation [44], and that pauses within this range can influence listener impressions [45]. We exclude turns shorter than 5 seconds and remove cases where the speech immediately before or after the pause consists of backchannel responses. For the smooth turn-taking task, we compute the gap between each pair of consecutive turns and retain only those where all gaps are less than 0.4 seconds. This threshold is based on findings that smooth conversational transitions in English typically occur within 200–250 milliseconds [46], and we conservatively set 0.4 seconds as an upper bound. We further require no overlapping speech, each turn to be longer than 4 seconds, and that the utterances are not backchanneling. To give evaluated models sufficient time to respond, we append 5 seconds of silence at the end of the input stream.
After automatic filtering, all retained segments are manually reviewed with dual-channel evidence (whether the partner is silent or entering, and whether there is hesitation or fluent continuation).
@@
Synthetic Data (User Interruption, Pause Handling): To address the scarcity of user interruptions in existing public datasets, we generated synthetic dialogues using GPT-4o [48], which include both contextual turns and interruption turns. Text-to-Speech synthesis was performed using ChatTTS [^6] https://github.com/2noise/ChatTTS, with 10 different speaker voices randomly assigned across samples to enhance diversity. For each input stream fed to the evaluated model, the interrupting speech is played around 7 seconds after the preceding utterance. Additionally, we append 15 seconds of silence after the interruption to allow the model time to respond.
@@ IV Models Under Evaluation
Since many models do not release complete speech-to-speech checkpoints (see Table I, column “S2S Release”), currently we evaluate models with a publicly available speech-to-speech model and inference pipeline. We also include the recently released Gemini Live API as a representative commercial model due to its user-friendly interface.
  • dGSLM [26]: A textless speech-to-speech model that generates natural conversations directly from audio. It uses (1) a HuBERT+k-means encoder [49], (2) a dual-tower Transformer for two-channel dialogue modeling, and (3) a HiFi-GAN decoder. Trained on 2,000 hours of phone calls [50], it captures both linguistic and paralinguistic cues. Originally non-interactive, we adapt it for live interaction using the official implementation [^7] https://github.com/facebookresearch/fairseq/tree/main/examples/textless_nlp/dgslm with our modifications.
  • Moshi [7]: A real-time speech-to-speech system combining a 7B LLM (Helium) and neural codec (Mimi) via residual vector quantization. It features an “Inner Monologue” step to improve fluency and supports overlapping speech and interruptions through a multi-stream architecture. We use the official implementation [^8] https://github.com/kyutai-labs/moshi.
@@ VI Conclusion
In this paper, we introduced Full-Duplex-Bench, a benchmark designed to evaluate critical aspects of full-duplex spoken dialogue models. Our framework targets key interaction dimensions—pause handling, backchanneling, smooth turn-taking, and user interruption management—addressing limitations of existing benchmarks that primarily focus on half-duplex settings or coarse corpus-level metrics. To ensure systematic and reproducible evaluation, we propose automatic metrics tailored to real-time interaction. Experiments on full-duplex models reveal distinct model features and highlight areas for improvement. By releasing our data and behavior-specific metrics, we hope Full-Duplex-Bench provides a practical foundation for evaluating full-duplex spoken dialogue systems.
VII Limitation and Future Work
Our framework does not yet link described behaviors to human preferences; users need to determine what constitutes desirable or undesirable behavior according to their specific goals. Future work can integrate human judgment studies to provide the preference. The present analysis is limited to English, and extending the framework to other languages will be essential for assessing cross-linguistic generality.